Sol Summary – November 29th

Sol 2

Title: An Intriguing Find…

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I set off to explore the area to the South of Kissing Camel Range and West of Cow Dung Rd. First I investigated some of the washes crossing the road to see what I could find. In the first wash that I stopped in, I found pieces of petrified wood amongst other very interesting specimens. I’ll definitely have to explore this drainage basin later.

From there, I proceeded further south, until I reached a large wash near White Rock Canyon. I only walked about 2 minutes up, when I found a very large specimen that looked like mineralized bone marrow. Compared to the basalt ejecta littering the wash, this specimen looked completely different. The holes were all aligned in the same direction, and were about the same size. The basalt pieces had gas bubbles, but were of random size, with no obvious pattern. The size of the marrow was very large, with 3-4 mm spacing. If this was indeed a piece of a dinosaur bone, then that meant there is a very large dinosaur fossil somewhere in that drainage basin.

I left the specimen where I found it, planning to come back and visit later. After spending over 2 hours in this area, I didn’t find much more, though there is still quite a bit of ground to cover. I returned to the hab by around 2:00 PM, and informed Shannon on what I had found. On our way to town, we investigated the specimen that I had found in the wash. Shannon wasn’t sure what it was, but we confirmed that there weren’t any other rocks in that area that looked like it.  She figured that if it was indeed a remnant of a dinosaur bone, that it would have to come from a very large dinosaur, much larger than a Utahraptor or velociraptor. We plan on surveying this area further to see what we can find.

Soon after, we ran into town, found a socket wrench kit that would work with the bolts on the new rovers, and organized the next supply of food for Crew 184. We returned to hab after finishing our business in town.

Look Ahead Plan:

Tomorrow and the day after, I will be extensively searching the drainage basin where I found that specimen for large fossil exposures. Since Shannon told me that she hasn’t extensively explored this area, what I may find there could be very interesting.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA:

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.

Sol Summary – November 28th

Crew 183+ Sol Summary Report 28November2017

Sol 1

Title: Departure of Crew 183

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

This was the last day for Crew 183. They packed the last of their belongings and headed off, back to Earth. After eating breakfast, Shannon and I started assembling the new rovers. It took over an hour to get just one close to being ready to go, as the rovers didn’t come with assembly instructions. Unfortunately, we don’t have a socket wrench that is compatible with the large bolts used with the rovers. Hopefully we’ll be able to find a socket large enough to work.

At around 2:30 PM, we headed off to Cactus Rd, to rediscover a road that was thought to be lost and forgotten to time. Our goal was to see if this area could be accessed and explored by future EVAs. It was obvious it hadn’t been used by ATVs or rovers in years, as the most prominent tracks were from cows. The cow tracks made the road extremely bumpy, slowing our progress significantly.

Once we reached the end, we found a wonderful place to access the Summerville Formation, an easy 100 meter walk. There were these red and yellow pieces of chert-like rocks scattered everywhere that could be petrified dinosaur poop.

Once we finished here, we made our way to the end of Yellow Zebra road and found an amazing place for future EVAs. This area could easily take several hours to explore adequately. As it got closer to dark, we made our way back to the Hab, and I resumed my soil analysis in the Lab.

Look Ahead Plan:

Tomorrow, we plan to gather food supplies for the next incoming crew. I also intend to start exploring the region south of Kissing Camel range, to see if I find anything interesting. There is also a road in the area we conducted our EVA today that we still need to explore, called Maxwell Montes Rd. We may either do this tomorrow or sometime before Saturday, when Crew 184 arrives.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: End of Cactus Rd & Yellow Zebra Rd

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 28th

Crew 183 + Operations Report 28NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC  92% (Before generator is run at night)

Diesel –  50 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –  not known gallons.

Water (trailer) –  0 gallons.

Water (static) –   300 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours: not recorded (but crews need to record this information)

Beginning charge: 100

Ending charge: 95

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: We began putting the new rovers together
today.  I decided that we would construct the cage, but not modify the
seat at this time.  Instead, we will just remove the back seat rest,
which was what the Mars 160 crew wanted to do originally but I
insisted on the modification.  This will still be done on these but at
a later date because it is much more secure.  Maybe when Yusuke is
here with his crew, since he modified them last time and will be able
to help me to do it.  In the meantime, crews will still be able to use
them.  They will require about a month of short trips before they are
able to go on long ones like Deimos.

Summary of internet:  not known

Summary of suits and radios:  Ashton has recommended that all the old
radios that are sketchy be taken out of service.  I agree.  Four new
radios and headsets will be added to the Hab before the next crew
arrives.  Old sketchy ones will be removed.

Summary of Hab operations: Two new CO alarms were delivered to the
Hab.  We will mount one each on the lower and upper deck before the
new crew arrives.

Summary of GreenHab operations:  Two new planters were added to the
GreenHab.  They are for an herb garden and a greens garden.  I am
prepping the soil.  I turned the heater down to 55 degrees F.  Today I
left the door open all day.  Used 6 gallons of water for plants and
planters.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support:  We have bolts
that have 10.8 and 8.8 on them.  What does that mean in turns of
socket size, if anything?  We don’t have a socket set that fits the
bolts on the rovers, so are trying to figure out if they are metric or
English system.  Any information appreciated.

Sol Summary – November 27th

Sol 5 Summary Report

November 27, 2017

Julielynn Wong, Crew 183 Mission Commander

Summary Title: Our final day of rapid prototyping humanitarian supplies at MDRS.

Mission Status: Nominal.

Sol Activity Summary:

Our simulation ended at 0700 MST today.

During our mission, we successfully 3D printed designs created remotely by Medical Makers Labeeba Nusrat, Nazanin Soghrati, Sachin Doshi, Shamsa Qaadri, Susan Lee, Winnie Xu, Zak Chaudhary and Ziyuan (Steve) Xu.

Today Crew 183 continued to rapidly prototype several life-saving and life-changing supplies on-site using our humanitarian 3D printing system (H3DPS-LM).

Our crew continued to use our cycling ergometer system to follow the exercise countermeasures schedule for astronauts onboard the ISS and to power our humanitarian 3D printing system. Power generation data was collected non-invasively during training that crew members would have undertaken if no experiment existed.

Our meals were prepared by the HSO.

To celebrate the end of our mission, we took photos of Crew 183 and the 3D4MD humanitarian 3D printing, recycling, and cycling ergometer systems inside the RAMM (Repair and Maintenance Module) which is slated for use for engineering research in 2018.

We are now packing up our equipment into our Tesla Model S electric vehicle for the return trip home.

Many thanks to Shannon, CAPCOM and the Mars Society team who made the inaugural Medical Makers MDRS mission possible.  We look forward to returning in the next field season.

 

Look Ahead Plan: Preparing to clean the hab and depart MDRS on Sol 6.

Anomalies in work:  Out of sim.

Weather:  Sunny with occasional winds.

Crew Physical Status:  Healthy

EVA: Crew Astronomer, Crew Engineer, Greenhab Officer, and HSO participated in an EVA led by our Crew Geologist.

 

Reports to be filed:

GreenHab Report

Operations Report

EVA Report

Astronomer Report

 

Support Requested: None

Operations Report – November 27th

EVA #5 Report

Today our EVA was conducted out-of-sim, with the purpose of further testing our 3D scanning and imagery technology out on the field. Jordan, Ashton, Toruwa, Dean, and I took Deimos and three ATVs to Lith Canyon. We left the hab at approximately 12:20 PM. On the way, we stopped by the location where we obtained the specimens and return them as we found them before.

After spending some time there, we arrived to the Lith Canyon turnout by around 1:15 PM. We walked to the ridge overlooking the canyon and conducted our activities in this area. We stayed here until about 2:30 PM, and made our way back to the Hab by 3:10 PM.

Have a good night Mission Support!

Brandon Ferguson

EVA Officer

GreenHab Report – November 27th

Sol 5 – 27/11/2017

Green Hab Officer Jordan Bowden

Average temperature: 28.5 C in Greenhouse / 10 C Outside

Time 11:30 16:30
Temperature inside (°C) 30 C 26 C
Temperature outside (°C) 8C 12 C
% Humidity 16% 16 %


Changes to crops:  Everything is growing well.

Daily water usage for crops: 3 Gallons

Time(s) of watering for crops:  11:00, 14:30
Daily watering and amount of water used:  2 gallons used at 11:30, 1 gallon used at 16:30. 3 gallons total.

Narrative: Greetings CapCom,

Unfortunately my time on Mars is coming to a close, and as such, today was my last full day taking care of the Green Hab. Things have been coming along steadily in the space, and I’m happy to report that all of the plants seem healthier and bigger than when I arrived; as a non-expert gardner, this is the best I could really hope for.

Today we went on our final EVA: the sky was cloudy and the winds very strong. This weather was totally different than the rest of our time here on Mars: on other EVAs, the sun was so strong that we were sweating in spite of the actual outside temperature. Overall, it was wonderful to see the scenery and get some good work done outside the Hab.

Lastly, I wanted to emphasize how much I have enjoyed our time here, and how transformative the impact this time has had been on our work and research. I’m amazed by this space and the community that keeps it going, and I can’t say enough how grateful I am for this experience.

Let me know if you need any electronics designed or parts 3D printed for the upcoming Aquaponics system in the Green Hab, I could send them out to Mars if it’s feasible.

You’re the best, CapCom. Thanks for your hard work.

Signing off,

Jordan Bowden
Green Hab Officer

Astronomy Report – November 26th

Astronomy/Science Officer Toluwa Fayemi
As far as our relatively uneventful stay on the Mars Desert Research Station goes, today was a bit more on the eventful side.
The morning started late and we jumped straight into preparing our experiments for the day. Jordan and I spent a good amount of time in the Geology Lab, testing our field 3D scanner on the specimen we retrieved from our EVA the previous day. The results were promising, but we expected to run into some issues in the field. Surely enough, it was difficult to scan an object 360º around while carrying the laptop with us without ‘losing’ the object in the scanner field, particularly in this terrain where the specimen of interest tend to be identical in colour to the ground around it. The sunlight also made it extremely difficult to see the laptop screen, making it difficult to monitor how well we were actually scanning the object.
During the same EVA, we managed to get some exciting photography/videography of the surrounding landscape which may be useful for future exploration/EVAs.
As far as astrophotography, I have been experimenting with me Canon t4i DSLR on various shutter speeds, apertures and ISO settings to continually improve my results. Some of the pictures and the settings used to obtain them are shared below.
This long exposure shot was amongst my first, set at f5.0 and an ISO of 128000, hence the grainy-looking ambient light around the hill. The shutter was open for 30 seconds.
This shot was taken with a briefer shutter, 25seconds and an f-stop of 6.3. This screenshot doesn’t quite do it justice but I’m pleased with the results of this settings modification. I believe the ISO was set to 64000 for this one.
This is the second shot I took with the HAB in view. I set the ISO to 64000 on this one, lower than the 128000 of the previous. The ambient light from the HAB was a tad too much for an ISO so high. This appears to be the better shot. Both shots had a 25 second shutter time and an f-stop of 6.3.